# Quantum Buffer Design Using Petri Nets

### v1

@@ A. Quantum S-R Flip-Flop Runs
We tested the design of Q-S-R flip-flop in Section 3 on IBM quantum computers under three different $SR$ settings, i.e., $|00\rangle$, $|10\rangle$ and $|01\rangle$ when the $QQ^{\prime}$-state is one of only two valid present states, i.e., $|10\rangle$ and $|01\rangle.$ As in a classical S-R flip-flop, the $|11\rangle$ input combination is not allowed for the Q-S-R flip-flop. We performed these runs by measuring both $Q$ and $Q^{\prime}$ as well as by measuring either $Q$ or $Q^{\prime}$ as depicted in Figure 15 and 16. Both runs were repeated 100 times for each test case. In two qubit measurements in Figure 15, the bars represent the frequencies of $|00\rangle,|01\rangle,|10\rangle,|11\rangle$ from left to right. For the set input, i.e., $|SR\rangle=|10\rangle,$ the Q-S-R flip-flop’s next state, $|QQ^{\prime}\rangle$ was measured $|10\rangle$ 30 and 29 times, respectively in the present states $|10\rangle$ and $|01\rangle,$ both more frequently than the other three states as shown in Figure 15(a) and 15(b). For the reset input, i.e., $|SR\rangle=|01\rangle,$ the next state $|QQ^{\prime}\rangle$ was measured $38$ and $39$ times, respectively in the present states $|10\rangle$ and $|01\rangle,$ again more frequently than the other states as shown in Figure 15(c) and 15(d). When $|SR\rangle$ is set to $|00\rangle$, the next state $|QQ^{\prime}\rangle$ was measured $|10\rangle$ and $|01\rangle$ $43$ and $38$ times, respectively in the present states $|10\rangle$ and $|01\rangle$ as shown in Figure 15(e) and 15(f). Similarly in one qubit measurement in Figure 16, we note that, for the set input, i.e., $|SR\rangle=|10\rangle,$ the Q-S-R flip-flop’s next state, $|Q\rangle=|1\rangle$ was measured $66$ and $76$ times, respectively in the present states $|10\rangle$ and $|01\rangle$ as shown in Figure 16(a) and 16(b). For the reset input, i.e., $|SR\rangle=|01\rangle,$ the next state $|Q\rangle=|0\rangle$ was measured $60$ and $55$ times, respectively in the present states $|10\rangle$ and $|01\rangle$ as shown in Figure 16(c) and 16(d). When $|SR\rangle$ is set to $|00\rangle$, the next state $|Q\rangle=|1\rangle$ was measured $62$ times, and $|Q\rangle=|1\rangle$ was measured $57$ times, respectively in the present states $|10\rangle$ and $|01\rangle$ as shown in Figure 16(e) and 16(f). These measurements suggest that reading both $Q$ and $Q^{\prime}$ leads to more measurement errors than reading only $Q.$
(a) $SR=10,QQ^{\prime}=10$
(b) $SR=10,QQ^{\prime}=01$
(c) $SR=01,QQ^{\prime}=10$
(d) $SR=01,QQ^{\prime}=01$
(e) $SR=00,QQ^{\prime}=10$
(f) $SR=00,QQ^{\prime}=01$
Figure 15: Measurement outcomes of $Q$ and $Q^{\prime}$ on QSR circuit using IBM Machine.
(a) $SR=10,QQ^{\prime}=10$
(b) $SR=10,QQ^{\prime}=01$
(c) $SR=01,QQ^{\prime}=10$
(d) $SR=01,QQ^{\prime}=01$
(e) $SR=00,QQ^{\prime}=10$
(f) $SR=00,QQ^{\prime}=01$
Figure 16: Measurement outcomes of qubit $Q$ on QSR circuit using IBM Machine.
B. SISO QPN Simulations
@@
We implement the SISO quantum buffer that was described in Figure 5 with three data $q$-tokens ($n=3$) $d_{1},d_{2},d_{3}$ in $P_{I}$ and two ancillary $q$-tokens ($m=2$) $z_{1},z_{2},$ in $P_{A}$ where $a_{1}=|10\rangle,a_{2}=|1\rangle,a_{3}=|1\rangle$ and $z_{1}=z_{2}=|0\rangle$ as shown in Figure 17(a) and 17(b). Initially $T_{1}$ is enabled and when its fires, it forwards the data $q$-token $d_{1}$ to $P_{O}$ and ancillary $q$-token $z_{1}$ to $P_{A1}$ as shown in Figure 17(c) and Figure 17(d). Again, $T_{1}$ is enabled and when its fires, it forwards the data $q$-token $d_{2}$ to $P_{O}$ and ancillary $q$-token $z_{2}$ to $P_{A1}$ as shown in Figure 17(g) and 17(i). The data $q$-token $d_{3}$ will remain in $P_{I}$ as there is no ancillary $q$-token left in $P_{A}$ to fire $T_{1}$ as shown in Figure 17(h). We note that to process two data $q$-tokens, two transitions are fired.
@@
C. SIMO QPN Simulations
We implement the SIMO quantum buffer that was described in Figure 9 with two output places $P_{O1},P_{O2}$, two transitions $T_{1},T_{2}$, four data $q$-tokens ($n=4$) $d_{1},d_{2},d_{3},d_{4}$ in $P_{I}$ and three ancillary $q$-tokens ($m=3$) $z_{1},z_{2},z_{3}$ in $P_{A}$ where $d_{1}=d_{3}=d_{4}=|1\rangle,d_{2}=|0\rangle$ and $z_{1}=z_{3}=|1\rangle,z_{2}=|0\rangle$ as shown in Figure 18 (a) and 18 (b). Initially, the first ancillary $q$-token $z_{1}=|1\rangle$, which enables the transition $T_{2}$ and when its fires, it forwards the data and ancillary $q$-token $d_{1}$ and $z_{1}$ to $P_{O2}$ and $P_{A1}$. The next ancillary $q$-token $z_{2}=|0\rangle$, which enables the transition $T_{1}$ and when its fires, it forwards the data and ancillary $q$-token $d_{2}$ and $z_{2}$ to $P_{O1}$ and $P_{A1}$. Lastly, the ancillary $q$-token $z_{1}=|1\rangle$ enables the transition $T_{2}$ and when its fires, it forwards the data and ancillary $q$-token $d_{3}$ and $z_{3}$ to $P_{O2}$ and $P_{A1}$ as shown in Figure 18(c) and 18(f), whereas data $q$-token $d_{4}$ will remain in $P_{I}$ as there is no ancillary $q$-token left in $P_{A}$ as shown in Figure 18(e).
@@ C. SIMO QPN Simulations
D. Priority QPN
We implement the priority quantum buffer that was described in Figure 13 with three data $q$-tokens ($n=3$) $d_{1},d_{2},d_{3}$ and four ancillary $q$-tokens ($m=2$) $z_{1},w_{2},z_{1},z_{2}$ where $d_{1}=|0\rangle,d_{2}=d_{3}=|1\rangle$ and $w_{1}=w_{2}=z_{1}=z_{2}=|0\rangle$. The data and ancillary $q$-tokens $d_{1}$ and $w_{1},w2$ are placed in low priority group $P_{I1}$ and $P_{A1}$ as shown in Figure 19 (a) and 19 (b), whereas data and ancillary $q$-tokens $d_{2},d_{3}$ and $z_{1},z_{2}$ are placed in high priority group $P_{I2}$ and $P_{A2}$ as shown in Figure 19(c) and 19(d). Initially, the enabled transitions are $T_{1}$ and $T_{2}$, where $T_{1}$ will fire once and $T_{2}$ will fire twice, when $T_{1}$ fires, it forwards the data and ancillary $q$-token $d_{1}$ and $w_{1}$ to $P_{DA1}$, when $T_{2}$ fires, it forwards the data and ancillary $q$-token $d_{2},d_{3}$ and $z_{1},z_{2}$ to $P_{DA2}$. If $P_{DA2}$ is empty, then $T_{3}$ will fire, otherwise $T_{4}$ fires. When $T_{3}$ fires, it forwards the data $q$-token $d_{1}$ to $P_{O}$ and ancillary $q$-token $w_{1}$ to $P_{A3}$, and when $T_{4}$ fires, it forwards the data $q$-token $d_{2},d_{3}$ to $P_{O}$ and ancillary $q$-token $z_{1},z_{2}$ to $P_{A3}$ as shown in Figure 19(e) and 19 (g), whereas ancillary $q$-token $w_{2}$ will remain in $P_{A}$ as there is no data $q$-token left in $P_{I1}$ as shown in Figure 19(f).

### v2

@@ A. Quantum S-R Flip-Flop Runs
We implemented the two Q-S-R flip-flop designs on IBM Brisbane, QPU-Eagle-r3, Version 1.1.41 using 6 qubits.
A.1. Quantum S-R Flip-Flop Circuit 1 (Figure 3)
The quantum diagram for Circuit 1 was created on the IBM composer as shown in Figure 16. The circuit was tested under possible $SR$ configurations, ($|00\rangle$, $|10\rangle$, and $|01\rangle$), with the $QQ^{\prime}$ state restricted to two valid present states, $|10\rangle$ and $|01\rangle$. Each test case was run 100 times, and the results are shown in Figures 17. The vertical axis represents the frequency of the output occurrences, while the horizontal axis corresponds to the sequence of output states: $SRQ^{\prime}QA_{1}A_{2},$ where $A_{1}$ and $A_{2}$ represent ancillary qubits. We analyzed both the combined and individual measurement outcomes such as SRQQ, QQ, and Q, to ensure a thorough evaluation of the circuit’s behavior under different input configurations as shown in Table 2. For the set input, i.e., $|SR\rangle=|10\rangle,$ the Q-S-R flip-flop’s combined and individual next states, $|SRQQ^{\prime}\rangle=|1010\rangle,|QQ^{\prime}\rangle=|10\rangle$ and $|Q\rangle=|1\rangle$ were measured $7,34,62$ and $6,28,53$ times, respectively in the present states $|QQ^{\prime}\rangle=|10\rangle$ and $|QQ^{\prime}\rangle=|01\rangle$. For the reset input, i.e., $|SR\rangle=|01\rangle,$ the combined and individual next states, $|SRQQ^{\prime}\rangle=|0101\rangle,|QQ^{\prime}\rangle=|01\rangle$ and $|Q=0\rangle$ were measured $5,30,57$ and $11,39,59$ times, respectively in the present states $|QQ^{\prime}\rangle=|10\rangle$ and $|QQ^{\prime}\rangle=|01\rangle$. When $|SR\rangle$ is set to $|00\rangle$, the combined and individual next states, $|SRQQ^{\prime}\rangle=|1010\rangle,|QQ^{\prime}\rangle=|10\rangle$ and $|Q=1\rangle$ were measured $8,34,52$ times in the present state, $|QQ^{\prime}\rangle=|10\rangle$ and $|SRQQ^{\prime}\rangle=|0101\rangle,QQ^{\prime}\rangle=|01\rangle$ and $|Q=0\rangle$ were measured $8,31,51$ in the present state $|QQ^{\prime}\rangle=|01\rangle$.
Figure 16: Quantum Implementation of Q-S-R Flip-Flop Circuit 1.
(a) Initial input $SR=|10\rangle,QQ^{\prime}=|10\rangle$
(b) Initial input $SR=|10\rangle,QQ^{\prime}=|01\rangle$
(c) Initial input $SR=|01\rangle,QQ^{\prime}=|10\rangle$
(d) Initial input $SR=|01\rangle,QQ^{\prime}=|01\rangle$
(e) Initial input $SR=|00\rangle,QQ^{\prime}=|10\rangle$
(f) Initial input $SR=|00\rangle,QQ^{\prime}=|01\rangle$
Figure 17: Measurement outcomes of Q-S-R Flip-Flop Circuit $1$
Input     Q Output Counts
S R Q Q     SRQQ QQ Q
$|1\rangle$ $|0\rangle$ $|1\rangle$ $|0\rangle$     $|1010\rangle,7$ $|10\rangle,34$ $|1\rangle,62$
$|1\rangle$ $|0\rangle$ $|0\rangle$ $|1\rangle$     $|1010\rangle,6$ $|10\rangle,28$ $|1\rangle,53$
$|0\rangle$ $|1\rangle$ $|1\rangle$ $|0\rangle$     $|0101\rangle,5$ $|01\rangle,30$ $|0\rangle,57$
$|0\rangle$ $|1\rangle$ $|0\rangle$ $|1\rangle$     $\,\,\,|0101\rangle,11$ $|01\rangle,39$ $|0\rangle,59$
$|0\rangle$ $|0\rangle$ $|1\rangle$ $|0\rangle$     $|0010\rangle,8$ $|10\rangle,34$ $|1\rangle,52$
$|0\rangle$ $|0\rangle$ $|0\rangle$ $|1\rangle$     $|0001\rangle,8$ $|01\rangle,31$ $|0\rangle,51$
Table 2: Compiled measurement outcomes of Q-S-R Flip-Flop Circuit $1$
A.2. Quantum S-R Flip-Flop Circuit 2 (Figure 4)
The quantum Circuit 2 was created on the IBM composer as shown in Figure 18. Like Circuit 1, Circuit 2 was also tested under possible $SR$ configurations, ($|00\rangle$, $|10\rangle$, and $|01\rangle$) with the $QQ^{\prime}$ state restricted to two valid present states, $|10\rangle$ and $|01\rangle$. Each test case was run 100 times, and the results are shown in Figures 19. The vertical axis represents the frequency of the output occurrences, while the horizontal axis corresponds to the sequence of output states: $SRA_{1}A_{2}Q^{\prime}Q,$ where $A_{1}$ and $A_{2}$ denote ancillary qubits as before. The circuit’s behavior under different input configurations is tabulated for different sets of output measurements in Table 3. For the set input, i.e., $|SR\rangle=|10\rangle,$ the Q-S-R flip-flop’s combined and individual next states, $|SRQQ^{\prime}\rangle=|1010\rangle,|QQ^{\prime}\rangle=|10\rangle$ and $|Q\rangle=|1\rangle$ were measured $63,71,85$ and $53,62,86$ times, respectively in the present states $|QQ^{\prime}\rangle=|10\rangle$ and $|QQ^{\prime}\rangle=|01\rangle$. For the reset input, i.e., $|SR\rangle=|01\rangle,$ the combined and individual next states, $|SRQQ^{\prime}\rangle=|0101\rangle,|QQ^{\prime}\rangle=|01\rangle$ and $|Q=0\rangle$ were measured $47,74,90$ and $63,75,90$ times, respectively in the present states $|QQ^{\prime}\rangle=|10\rangle$ and $|QQ^{\prime}\rangle=|01\rangle$. When $|SR\rangle$ is set to $|00\rangle$, the combined and individual next states, $|SRQQ^{\prime}\rangle=|1010\rangle,|QQ^{\prime}\rangle=|10\rangle$ and $|Q=1\rangle$ were measured $49,57,77$ times in the present states $|QQ^{\prime}\rangle=|10\rangle$ and $|SRQQ^{\prime}\rangle=|0101\rangle,QQ^{\prime}\rangle=|01\rangle$ and $|Q=0\rangle$ were measured $57,71,79$ in the present states $|QQ^{\prime}\rangle=|01\rangle$.
Figure 18: Circuit $2$ implementation of a Q-S-R flip-flop.
(a) Initial input $SR=|10\rangle,Q=|1\rangle$
(b) Intial input $SR=|10\rangle,Q=|0\rangle$
(c) Initial input $SR=|01\rangle,Q=|1\rangle$
(d) Initial input $SR=|01\rangle,Q=|0\rangle$
(e) Initial input $SR=|00\rangle,Q=|1\rangle$
(f) Initial input $SR=|00\rangle,Q=|0\rangle$
Figure 19: Circuit 2 measurement outcomes.
Input     Output
S R Q     SRQQ QQ Q
$|1\rangle$ $|0\rangle$ $|1\rangle$     $|1010\rangle=63$ $|10\rangle=71$ $|1\rangle=85$
$|1\rangle$ $|0\rangle$ $|0\rangle$     $|1010\rangle=53$ $|10\rangle=62$ $|1\rangle=86$
$|0\rangle$ $|1\rangle$ $|1\rangle$     $|0101\rangle=47$ $|01\rangle=74$ $|0\rangle=90$
$|0\rangle$ $|1\rangle$ $|0\rangle$     $|0101\rangle=63$ $|01\rangle=75$ $|0\rangle=90$
$|0\rangle$ $|0\rangle$ $|1\rangle$     $|0010\rangle=49$ $|10\rangle=57$ $|1\rangle=77$
$|0\rangle$ $|0\rangle$ $|0\rangle$     $|0001\rangle=57$ $|01\rangle=71$ $|0\rangle=79$
Table 3: Compiled measurement outcomes of Q-S-R flip-flop Circuit $2$
A.3. Quantum S-R Flip-Flop Via A Quantum J-K Flip-Flop
Figure 20: Implementation of a Q-J-K flip-flop circuit [31].
(a) Initial input $JK=|10\rangle,Q=|1\rangle$
(b) Intial input $JK=|10\rangle,Q=|0\rangle$
(c) Initial input $JK=|01\rangle,Q=|1\rangle$
(d) Initial input $JK=|01\rangle,Q=|0\rangle$
(e) Initial input $JK=|00\rangle,Q=|1\rangle$
(f) Initial input $JK=|00\rangle,Q=|0\rangle$
Figure 21: Q-J-K circuit measurement outcomes.
As stated in Section 3, a quantum S-R flip-flop can be obtained from a quantum J-K flip-flop with $(J=|0\rangle,K=|0\rangle),(J=|0\rangle,K=|1\rangle),(J=|1\rangle,K=|0\rangle)$ settings, and avoiding the $(J=|1\rangle,K=|1\rangle)$ setting. To compare the performance of the two Q-S-R flip-flop circuits to the performance of a quantum J-K flip-flop, we implemented the quantum J-K flip-flop that was reported in [31] on the IBM Brisbane, QPU-Eagle-R3, Version 1.1.41 as shown in Figure 20.
Input     Output
J K Q     QQ Q
$|1\rangle$ $|0\rangle$ $|1\rangle$     $|10\rangle=34$ $|1\rangle=51$
$|1\rangle$ $|0\rangle$ $|0\rangle$     $|10\rangle=42$ $|1\rangle=50$
$|0\rangle$ $|1\rangle$ $|1\rangle$     $|01\rangle=35$ $|0\rangle=44$
$|0\rangle$ $|1\rangle$ $|0\rangle$     $|01\rangle=45$ $|0\rangle=56$
$|0\rangle$ $|0\rangle$ $|1\rangle$     $|10\rangle=22$ $|1\rangle=43$
$|0\rangle$ $|0\rangle$ $|0\rangle$     $|01\rangle=31$ $|0\rangle=44$
Table 4: Compiled measurement outcomes of Q-J-K Flip-Flop.
Table 4 summarizes the measurements for $QQ^{\prime}$ and $Q$ outputs for possible $JK$ configurations, ($|00\rangle$, $|10\rangle$, and $|01\rangle$) with the $QQ^{\prime}$ state restricted to two valid present states, $|10\rangle$ and $|01\rangle.$ The results indicate that, in terms of valid next states measurements, Circuit 2 implementation of the Q-S-R flip-flop outperforms both Circuit 1 implementation of the Q-S-R flip-flop and the implementation of Q-J-K flip-flop in [31] by a significant margin while Circuit 1 implementation of the Q-S-R flip-flop and the implementation of Q-J-K flip-flop have comparable correct measurements of $QQ^{\prime}$ and $Q$ states. We also tabulated the quantum gate counts of the three quantum flip-flop implementations as well as the Q-J-K flip-flop design that was described in [30] in Table 5. The implementation of this Q-J-K flip-flop was not carried out as it uses a CV+ gate that is not directly available in IBM Composer and OpenQasm 2.0. Overall, Circuit 2 implementation of S-R-Q flip-flop outperforms the other quantum flip-flop implementations in terms of the validity of its output states and cost.
    NOT CV CV+ C-NOT CC-NOT CCC-NOT
Quantum J-K [30]     -- $2$ $1$ $8$- --
Quantum J-K [31]     $6$- -- $14$ $4$-
Circuit-1     $6$- -- $2$ $1$ $12$
Circuit-2     $8$- -- $1$ $3$-
Table 5: Gate counts of Q-S-R and Q-J-K flip-flops.
B. Quantum Buffer Simulations
We simulated all quantum buffer designs on Jupyter [34] in Python and IBM Qiskit [32].
B.1. The SISO QPN Simulations
@@
We implemented the SISO quantum buffer that was described in Figure 6 with three data $q$-tokens ($n=3$) $d_{1},d_{2},d_{3}$ in $P_{I}$ and two ancillary $q$-tokens ($m=2$) $z_{1},z_{2},$ in $P_{A}$ where $a_{1}=|10\rangle,a_{2}=|1\rangle,a_{3}=|1\rangle$ and $z_{1}=z_{2}=|0\rangle$ as shown in Figure 22(a) and 22(b). Initially, $T_{1}$ is enabled and when its fires, it forwards the data $q$-token $d_{1}$ to $P_{O}$ and ancillary $q$-token $z_{1}$ to $P_{A1}$ as shown in Figure 22(c) and Figure 22(d). Again, $T_{1}$ is enabled and when its fires, it forwards the data $q$-token $d_{2}$ to $P_{O}$ and ancillary $q$-token $z_{2}$ to $P_{A1}$ as shown in Figure 22(g) and 22(i). The data $q$-token $d_{3}$ will remain in $P_{I}$ as there is no ancillary $q$-token left in $P_{A}$ to fire $T_{1}$ as shown in Figure 22(h). We note that to process two data $q$-tokens, two transitions are fired.
@@
B2. The SIMO QPN Simulations
We implemented the SIMO quantum buffer that was described in Figure 10 with two output places $P_{O1},P_{O2}$, two transitions $T_{1},T_{2}$, four data $q$-tokens ($n=4$) $d_{1},d_{2},d_{3},d_{4}$ in $P_{I}$ and three ancillary $q$-tokens ($m=3$) $z_{1},z_{2},z_{3}$ in $P_{A}$ where $d_{1}=d_{3}=d_{4}=|1\rangle,d_{2}=|0\rangle$ and $z_{1}=z_{3}=|1\rangle,z_{2}=|0\rangle$ as shown in Figure 23(a) and 23(b). Initially, the first ancillary $q$-token $z_{1}=|1\rangle$, which enables the transition $T_{2}$ and when its fires, it forwards the data and ancillary $q$-token $d_{1}$ and $z_{1}$ to $P_{O2}$ and $P_{A1}$. The next ancillary $q$-token $z_{2}=|0\rangle$, which enables the transition $T_{1}$ and when its fires, it forwards the data and ancillary $q$-token $d_{2}$ and $z_{2}$ to $P_{O1}$ and $P_{A1}$. Lastly, the ancillary $q$-token $z_{1}=|1\rangle$, which enables the transition $T_{2}$ and when its fires, it forwards the data and ancillary $q$-token $d_{3}$ and $z_{3}$ to $P_{O2}$ and $P_{A1}$ as shown in Figure 23(c) and 23(f), whereas data $q$-token $d_{4}$ will remain in $P_{I}$ as there is no ancillary $q$-token left in $P_{A}$ as shown in Figure 23(e).
@@
B3. The MISO QPN Simulations
We implemented the MISO quantum buffer that was described in Figure 8 with two input places $P_{I1},P_{I2}$, two transitions $T_{1},T_{2}$, three data $q$-tokens ($n=3$) $d_{1},d_{2}$ in $P_{I1}$ ,$d_{3}$ in $P_{I2}$ and two ancillary $q$-tokens ($m=3$) $z_{1},z_{2},z_{3}$ in $P_{A}$ where $d_{1}=d_{3}=|1\rangle,d_{2}=|0\rangle$ and $z_{1}=|1\rangle,z_{2}=|0\rangle$ as shown in Figure 24(a),  24(b) and 24(c). Initially, the first ancillary $q$-token $z_{1}=|1\rangle$, which enables the transition $T_{2}$ and when its fires, it forwards the data and ancillary $q$-token $d_{3}$ and $z_{1}$ to $P_{O}$ and $P_{A1}$. The next ancillary $q$-token $z_{2}=|0\rangle$, which enables the transition $T_{1}$ and when its fires, it forwards the data and ancillary $q$-token $d_{1}$ and $z_{2}$ to $P_{O}$ and $P_{A1}$ as shown in Figure 24(e) and 23(f), whereas data $q$-token $d_{2}$ will remain in $P_{I1}$ as shown in Figure 8(d).
(a) $d_{1},d_{2}$ in $P_{I1}$
(b) $d_{3}$ in $P_{I2}$
(c) $z_{1},z_{2}$ in $P_{A}$
(d) $d_{2}$ in $P_{I1}$
(e) $d_{3},d_{1}$ in $P_{O}$
(f) $z_{1},z_{2}$ in $P_{A1}$
Figure 24: A partial flow of data and ancillary $q$-tokens in a MISO quantum buffer example.
B4. The MIMO QPN Simulations
We implemented the MIMO quantum buffer that was described in Figure 12 with two input places $P_{I1},P_{I2}$, four transitions $T_{1},T_{2},T_{3},T_{4}$, three data $q$-tokens ($n=4$) $d_{1},d_{2}$ in $P_{I1}$ ,$d_{3},d_{4}$ in $P_{I2}$ and four ancillary $q$-tokens ($m=2$) $w_{1},w_{2},z_{1},z_{2}$ where $d_{1}=d_{3}=d_{4}|1\rangle,d_{2}=|0\rangle$ and $w_{1}=z_{1}=z_{2}=|1\rangle,w_{2}=|0\rangle$ as shown in Figure 25(a),  25(b),  25(c) and 25(d). Initially, the first ancillary $q$-token $w_{1}=|1\rangle$, which enables the transition $T_{2}$ and when its fires, it forwards the data and ancillary $q$-token $d_{3}$ and $w_{1}$ to $P_{DA}$, whereas the second ancillary $q$-token $w_{2}=|0\rangle$, which enables the transition $T_{1}$ and when its fires, it forwards the data and ancillary $q$-token $d_{1}$ and $w_{2}$ to $P_{DA}$ as shown in Figure 25(e) . The ancillary $q$-tokens in $P_{A1}$ $z_{1}=z_{2}=|1\rangle$, which enables the transition $T_{4}$ and when its fires, it forwards the data and ancillary $q$-token $d_{3},d_{1}$ and $w_{1},w_{2},z_{1},z_{2}$ to $P_{O}$ and $P_{A2}$ as shown in Figure 25(h) and 25(i), whereas data $q$-token $d_{2}$ and $d_{4}$ will remain in $P_{I1}$ and $P_{I2}$ as shown in Figure 25(f),  25(g) .
(a) $d_{1},d_{2}$ in $P_{I1}$
(b) $d_{3},d_{4}$ in $P_{I2}$
(c) $w_{1},w_{2}$ in $P_{A}$
(d) $z_{1},z_{2}$ in $P_{A1}$
(e) $w1,w2,d_{3},d_{1}$ in $P_{DA}$
(f) $d_{2}$ in $P_{I1}$
(g) $d_{4}$ in $P_{I2}$
(h) $d_{3},d_{1}$ in $P_{O2}$
(i) $w_{1},w_{2},z_{1},z_{2}$ in $P_{A2}$
Figure 25: A partial flow of data and ancillary $q$-tokens in a MIMO quantum buffer example.
B5. Priority QPN
We implement the priority quantum buffer that was described in Figure 14 with three data $q$-tokens ($n=3$) $d_{1},d_{2},d_{3}$ and four ancillary $q$-tokens ($m=2$) $w_{1},w_{2},z_{1},z_{2}$ where $d_{1}=|0\rangle,d_{2}=d_{3}=|1\rangle$ and $w_{1}=w_{2}=z_{1}=z_{2}=|0\rangle$. The data and ancillary $q$-tokens $d_{1}$ and $w_{1},w2$ are placed in low priority group $P_{I1}$ and $P_{A}$ as shown in Figure 26 (a) and 26 (b), whereas data and ancillary $q$-tokens $d_{2},d_{3}$ and $z_{1},z_{2}$ are placed in high priority group $P_{I2}$ and $P_{A1}$ as shown in Figure 26(c) and 26(d). Initially, the enabled transitions are $T_{1}$ and $T_{2}$, where $T_{1}$ will fire once and $T_{2}$ will fire twice, when $T_{1}$ fires, it forwards the data and ancillary $q$-token $d_{1}$ and $w_{1}$ to $P_{DA1}$, when $T_{2}$ fires, it forwards the data and ancillary $q$-token $d_{2},d_{3}$ and $z_{1},z_{2}$ to $P_{DA2}$. If $P_{DA2}$ is empty, then $T_{3}$ will fire, otherwise $T_{4}$ fires. When $T_{3}$ fires, it forwards the data $q$-token $d_{1}$ to $P_{O}$ and ancillary $q$-token $w_{1}$ to $P_{A2}$, and when $T_{4}$ fires, it forwards the data $q$-token $d_{2},d_{3}$ to $P_{O}$ and ancillary $q$-token $z_{1},z_{2}$ to $P_{A2}$ as shown in Figure 26(e) and 26 (g), whereas ancillary $q$-token $w_{2}$ will remain in $P_{A}$ as there is no data $q$-token left in $P_{I1}$ as shown in Figure 26(f).